A farmer sold bushels of wheat and bushels of oats. For 425 bushels he received $481.58. How many bushels of each did he sell, if he received $2.08 per bushel of wheat and $.94 per bushel of oats?
HELP!!! I need help organizing this information in a chart and showing my work. I have no idea...Thanks.

Answer by checkley77(12844)

2.08x+.94(425-x)=481.58
2.08x+399.50-.94x=481.58
1.14x=481.58-399.50
1.14x=82.08
x=82.08/1.14
x=72 bushels of wheat were sold.
425-72=353 bushels of oats were sold.
Proof:
2.08*72+.94*353=481.58
149.76+331.82=481.58
481.58=481.58